# Hiding Functions within Functions: Steganography by Implicit Neural Representations

### v1

@@ Abstract
Deep steganography utilizes the powerful capabilities of deep neural networks to embed and extract messages, but its reliance on an additional message extractor limits its practical use due to the added suspicion it can raise from steganalyzers. To address this problem, we propose StegaINR, which utilizes Implicit Neural Representation (INR) to implement steganography. StegaINR embeds a secret function into a stego function, which serves as both the message extractor and the stego media for secure transmission on a public channel. Recipients need only use a shared key to recover the secret function from the stego function, allowing them to obtain the secret message. Our approach makes use of continuous functions, enabling it to handle various types of messages. To our knowledge, this is the first work to introduce INR into steganography. We performed evaluations on image and climate data to test our method in different deployment contexts.
Index Terms:
Steganography, Implicit Neural Representation, Continuous Function
I Introduction
@@
Steganography is a technique used for secret communication, where confidential information is hidden within a seemingly innocuous cover data and transmitted through a public channel [1]. The cover data can take various forms, such as digital text, audio, images, videos, or 3D data. Traditional steganographic methods primarily focus on designing algorithms to subtly modify [3] or select [2] the cover media in order to embed the secret information. However, it is challenging to achieve a good balance between concealment [4][5] and embedding capacity.
@@
Recently, deep neural network-based steganography schemes have been proposed, which generally exhibit superior performance in comparison to traditional approaches. In deep steganography, the sender of the message employs an encoding network to embed confidential information into the cover medium, while the receiver of the message makes use of a message decoder to recover the secret information from the cover medium. However, deploying these deep steganography methods in practical applications poses two critical issues. Firstly, these methods necessitate the transmission of a large decoder to the message recipient, but it is insecure to transmit the secret decoder for another transmission. Secondly, a more concealed and serious problem arises when the decoder is present at the recipient’s end, which may arouse suspicion from analysts regarding the steganographic behavior.
Figure 1: We introduce the new problem of INR steganography: hiding function into function. Our proposed framework, StegaINR, can embed and recover multimodal data through INR to implement secure covert communication.
One possible solution is to treat the message decoder as ordinary binary data and hide it in a popular cover data such as an image or video using existing steganography schemes. However, due to the relatively large size of the message decoder, traditional steganography schemes require a significant amount of cover data to meet the limited capacity, which imposes a significant communication burden. Another solution [49] [50] is to embed the decoder network into another deep neural network that performs other tasks, such as image classification or segmentation. However, this stego network is even larger than the message extractor and also causes a communication burden. In addition to the concealed transmission of the message decoder, both of these solutions require the transmission of the stego data.
@@
To address the covert issue of message decoders, we propose a novel steganographic scheme based on implicit neural representation. Implicit neural representations (INR) parameterize the signal as a continuous function. In our scheme, as shown in Fig. 1, (1) we first represent the secret message as a continuous secret function. (2) By using a shared key, we extend the secret function (network) to construct a new network structure. (3) With fixed parameters of the secret function, we parameterize cover data using the extended function to obtain a stego function. The stego function serves as both an extractor and a stego data, enabling secure transmission over a public channel. (4) The recipient only needs to utilize the shared key to recover the secret message from the stego function. Other users without the key can only obtain the cover data from the stego function. Our proposed implicit representation-based steganographic technique has the following advantages:
@@
Capacity: Since implicit neural representations parameterize the secret message as a continuous function, the memory required for parameterized signals is independent of spatial resolution but only dependent on the complexity of the underlying signal. This allows us to achieve high-capacity steganography using continuous functions.
Communication Cost: In our method, the stego function itself serves as the cover data. We only need to transmit the stego function over a public channel to achieve covert communication. Furthermore, implicit representations of continuous functions typically have a smaller size. Compared to the literature [49] [50], this significantly reduces the communication overhead.
@@
Efficiency: Unlike [49] [50], our approach only requires parameterization on a single data sample, without the need for training on large-scale datasets for other tasks. Furthermore, training implicit representations requires transforming grid data into coordinate and feature-based representation. This representation enables us to randomly sample a subset of elements instead of representing all data, for example, selecting a certain number of pixels from the entire image. Training on these subsets not only eliminates direct dependence on data resolution but also saves ample memory and improves model training efficiency.
In particular, the recent development of neural implicit representation technology has made it a trend to widely distribute visual data by sharing implicit representation model weights. An example of this is the NeRF[8] for 3D representation. We believe that in the future, people will share various multimedia data contents that they capture as function formats, just as they currently share 2D images and videos online. In this context, we are also interested in the following research questions:1. It is common to inject information into 2D images or 3D content for steganography or ownership recognition. But when people share information through implicit representation functions, can we achieve information hiding? 2. How to define steganographic capacity when implementing embedding operations in network with implicit representation?
@@
II Related Work
II-A Deep Steganography
@@
Deep steganography refers to a technique that employs deep neural networks to implement covert communication. These methods are mainly divided into two categories. One category is based on the encoder-decoder networks. They unitize the powerful encoding ability of deep neural networks to encode the cover data and message into the stego data. Baluja[59] first proposed the use of deep neural networks as encoders to hide secret images into other images, and they also trained a decoder to recover the message. Zhu et al. [9] introduced a noise layer into this framework, allowing the decoder to achieve robust message recovery. Jing et al. [61] unified the encoder and decoder into a reversible network to achieve deep steganography. They also [10] implemented multi-image embedding to achieve high-capacity steganography. Xu et al. [11] proposed a flow-based robust reversible image steganography framework. Recently, a special method for steganography through fixed neural networks [12] has been proposed, which adds controllable noise to the carrier to avoid training the message extraction network. These methods require an existing natural and realistic cover to construct a stego data.
@@
Another category of methods uses deep generative models, such as GANs, diffusion models, flow models, and autoregressive encoders to construct cover data, stego data or modification policies. Volkhonskiy et al.[14] first used GANs to generate an original cover image to counter steganalysis. Liu et al. [15] used a category-controlled GAN to establish a mapping relationship between the message and generated images to hide information. Similar to [14], Hu et al. [16] establishes a mapping relationship between messages and noise images and then uses the GAN to generate a stego image directly. Zhang et al. [17] regarded the generator as an image sampler and extracted the message using a fixed message decoder called the Carden Grille, Tang et al. [18] [19] used GANs to generate distortion cost functions to implement cover modification steganography. Chen et al. [20] directly generates stego using variational autoencoders [31] [42] (VAE) and flow model-based generators [33]. Yang et al. [34] proposed a new framework for generative steganography using autoregressive models, which uses the pixel CNN+ to generate hidden steganography. Zhang et al. [35] also used autoregressive models to design steganography schemes and implemented pixel-level information hiding using Pixel CNN+ and arithmetic coding. Liu et al. [36] decoupled images into structural and textural features using autoencoders, enabling stable feature extraction from structural features to improve the accuracy of secret message extraction. Recently, there have also been schemes using diffusion models[37][38] to construct steganography. Wei et al. [39] embed the message using modified DC coefficients of the diffusion noise, and then a reverse diffusion model restores the noise to a natural image. Finally, the message is extracted using a diffusion process and IDCT transform. Similar to Hu et al. [16], Kim et al.[40] establishes a mapping relationship between messages and noise images and then used a diffusion model to generate stego data. Yu[41] uses the image conversion ability of the diffusion model and the robustness of noise to adopt a stable diffusion conditional model to implement an image steganography.
@@
In order to achieve high-quality coding of cover data and decoding of messages, these deep learning-based steganography schemes usually require a lot of computing resources to obtain powerful encoders (generator) and decoders. These networks are usually trained on grid-like data such as images, making the size of these encoders and decoders directly related to the resolution of the underlying grid data. Except for Zhang et al. [17] and FNNS [12], all the above methods need to transfer the message extractor with a large amount of data to the message receiver, which not only increases the communication burden but also increases the risk of exposing steganographic behavior. To solve this problem, Li et al. [49] [50] disguise a steganography network (called a secret DNN model)by ordinary deep neural networks (called stego DNN) with image classification or segmentation tasks. In [49], they disguised the entire secret message extractor by selecting and adjusting some of the filters to retain their functionality in the secret task. In [50], they proposed a gradient-based filter integration scheme to insert interference filters into important positions in the secret DNN model to form a stego DNN model. Although these two methods ensure the security of the message decoder, they still do not avoid the problem of high communication burden. These two methods use image processing behavior (classification or segmentation) to conceal steganography behavior. They only provide a concealment strategy for message extractors in steganographic schemes. In this paper, we present a complete steganographic framework for multimedia data and other data, where we represent the message as a continuous function and embed this function into another continuous function. This not only provides security for message extractors but also eliminates the need for additional transmission of stego data, as the stego function itself acts as both a stego and extractor. Additionally, the resolution of the neural implicit representation is decoupled from the training data, and implicit model representations typically have smaller data sizes, thus avoiding high-load communication.
II-B Model Hiding
@@
With the widespread use of model data in the cloud, researchers have also paid extensive attention to information hiding in model data. Model hiding can be mainly divided into two aspects: Neural Network Steganography and Neural Network Watermarking. Neural network watermarking can be classified into three types: white-box watermarking, black-box watermarking, and no-box watermarking. In white-box watermarking, The verifier can access the internal structure of the network, including the weights, to verify the copyright. Uchida [42] proposed a method to embed a watermark consisting of 0s and 1s into the weights of the network through regularization, and the watermark can be extracted using a specific algorithm to verify the copyright. Black-box watermarking is suitable when the verifier cannot access the internal structure of the network and can only use remote API calls. A common approach [43] is to select specific samples as triggers and fine-tune the network to fit these triggers for copyright verification. No-box watermarking [44] is mainly used for copyright verification of generative networks. The strategy is to train the network to generate images that contain the watermark, and the verifier can directly verify the copyright from the generated images.
@@
Some scholars have proposed the model steganography where the model data is used as a stego data for hidden information. Yang et al. [45] embedded secret data into the convolutional layers of a given neural network (the cover network) during the network training process. Matrix multiplication is used to encode the parameters of the convolutional layers for data extraction. Chen et al. [46] used a deep neural network (DNN) to establish a probability density model of cover images and hide a secret image at a specific location of the learned distribution. They used SingGAN [47] to learn the distribution of patches in a cover image. During the patch distribution learning process, a deterministic mapping is fitted from a fixed set of noise maps (generated by embedding keys) to hide the secret image. Yang et al. [48] proposed a multi-source data hiding scheme for neural networks, where multiple senders can simultaneously send different secret data to a receiver using the same neural network. They achieved data embedding during the training process of the neural network, replacing post-training modifications to the network.
@@
II-C Information Hiding For NeRF
Recently, a special type of neural implicit representation for 3D data called Neural Radiance Fields (NeRF)[8] has received extensive research attention. Li et al. [51] proposed StegaNeRF, a method to embed hidden information in NeRF rendering. They designed an optimization framework that first trains a regular NeRF model and then performs a second training to achieve the rendering of the secret image while ensuring the original visual quality. Similarly, Luo et al. [52] proposed an anti-distortion rendering scheme by replacing the original color representation in NeRF with a watermarked color representation, ensuring stable extraction of the watermark information from the 2D images rendered by NeRF. Chen et al. [53] utilized the synthesized novel views by NeRF as an important basis for copyright verification and trained a watermark verification model using a parameterized approach. Similar to [53], Dong et al. [21] applied this technique to steganography and developed a Nerf-based steganography scheme. These approaches treat the neural radiance field as a new-view image generator, embedding secret information in the images from the new views. Therefore, an additional decoder is still required for message extraction. In this paper, we embeds the implicit representation function of the secret message into the implicit representation function of the cover data. We can directly extract the secret message from the stego function using the shared key between the sender and receiver, avoiding the need for the message extractor.
III Representing Data As Functions
In this section, we will provide a comprehensive overview of implicit neural representations, with a focus on using images as an illustrative example.
III-A Representing A Single Image With A Function
Suppose we have an image where $I[x;y]$ corresponds to the RGB value at pixel position $(x;y)$. We want to represent this image as a function $f:{{R}^{2}}\rightarrow{{R}^{3}}$, where $f(x;y)=(r;g;b)$ returns the RGB value at pixel position $(x;y)$. To achieve this, we parameterize the function $f_{\theta}$ with a multi-layer perceptron (MLP) with weighted parameters $\theta$, which is often referred to as an $implicit$ $neural$ $representation$. We can then learn this representation by minimizing the following loss function [29]:

### v2

@@
Abstract: Deep steganography utilizes the powerful capabilities of deep neural networks to embed and extract messages, but its reliance on an additional message extractor limits its practical use due to the added suspicion it can raise from steganalyzers. To address this problem, we propose StegaINR, which utilizes Implicit Neural Representation (INR) to implement steganography. StegaINR embeds a secret function into a stego function, which serves as both the message extractor and the stego medium for secure transmission on a public channel. Recipients only need to use a shared key to recover the secret function from the stego function, allowing them to obtain the secret message. Our approach employs continuous functions, enabling it to handle various types of messages. To our knowledge, this is the first work to introduce INR into steganography. We performed evaluations on image, climate data and NeRF synthetic dataset to test our method in different deployment contexts.
Keywords: Steganography; Implicit Neural Representation; Continuous Function
$\bullet$ Peng Luo is with the School of Cybersecurity, Northwestern Polytechnical University, Xi’an 710072, China and the School of Cryptography Engineering, Engineering University of PAP, Xi’an 710086, China. E-mail: lp_nwpu@mail.nwpu.edu.cn
$\bullet$ Jia Liu, Yan Ke, Qian Dang, Minqing Zhang are with the School of Cryptography Engineering, Engineering University of PAP, Xi’an 710086, China. Email: liujia1022@gmail.com; 15114873390@163.com; dangqian729@163.com; api_zmq@ 126.com.
$\bullet$ Dejun Mu are with the school of cybersecurity, northwestern polytechnical university, Xi’an 710072, China. Email: mudejun@mail.nwpu.edu.cn.
[…]
1 Introduction
@@
Steganography is a technique used for secret communication, where confidential information is hidden within a seemingly innocuous cover data and transmitted through a public channel[1]. The cover data can take various forms, such as digital text, audio, images, videos, or 3D data. Traditional steganographic methods primarily focus on designing algorithms respectively to subtly modify or select the cover media in order to embed the secret information[3, 2]. However, it is challenging to achieve a good balance between concealment[4, 5, 6] and embedding capacity.
Fig.  1: INR-based Steganography: Embedding a Secret Function into a Stego Function. StegaINR enables secure covert communication by embedding and recovering multimodal data via implicit neural representations.
@@
Recently, deep neural network-based steganography schemes have been proposed, which generally exhibit superior performance in comparison to traditional approaches. In deep steganography, the sender of the message employs an encoding network to embed confidential information into the cover medium, while the receiver of the message makes use of a message decoder to recover the secret information from the cover medium. However, deploying these deep steganography methods in practical applications poses two critical issues. Firstly, these methods necessitate the transmission of a large decoder to the message recipient, but it is insecure to transmit the secret decoder for another transmission. Secondly, a more concealed and serious problem arises when the decoder is present at the recipient’s end, which may arouse suspicion from analysts regarding the steganographic behavior. One possible solution is to treat the message decoder as ordinary binary data and hide it in a popular cover data such as an image or video using existing steganography schemes. However, due to the relatively large size of the message decoder, traditional steganography schemes require a significant amount of cover data to hide secret information, which imposes a significant communication burden. Another solution[7, 8] is to embed the decoder network into another deep neural network that performs other tasks, such as image classification or segmentation. However, this stego network is even larger than the message extractor and also causes a communication burden. In addition to the concealed transmission of the message decoder, both of these solutions require the transmission of the stego data.
@@
To address the covert issue of message decoders, we propose a novel steganographic scheme based on INR, which parameterizes the signal as a continuous function. In our scheme, as shown in Fig. 1, (1) we first represent the secret message as a continuous secret function, (2) by using a shared key, we extend the secret function (network) to construct a new network structure, (3) with fixed parameters of the secret function, we parameterize cover data using the extended function to obtain a stego function. The stego function serves as both an extractor and stego data, enabling secure transmission over a public channel. (4) The recipient only needs to utilize the shared key to recover the secret message from the stego function. Other users without the key can only obtain the cover data from the stego function. Our proposed INR based steganographic technique has the following advantages:
@@
Capacity: Since INR parameterize the secret message as a continuous function, the memory required for parameterized signals is independent of spatial resolution but only dependent on the complexity of the underlying signal. This allows us to achieve high capacity steganography using continuous functions.
Communication Cost: In the actual covert communication process, each transmitted secret message is different. In the scheme proposed in this paper, the secret function is INR of the secret message, and transmitting the secret function is essentially transmitting the secret message. In our method, the extended secret function (stego function) itself serves as the cover data. We only need to transmit the stego function over a public channel to achieve covert communication. Furthermore, implicit representations of continuous functions typically have a smaller size. Compared to [7, 8], this significantly reduces the communication cost.
@@
Efficiency: Unlike [7, 8], our approach only requires parameterization on a single data sample, without the need for training on large-scale datasets for other tasks. Furthermore, training implicit representations requires transforming grid data into coordinate and feature-based representation. This representation enables us to randomly sample a subset of elements instead of representing all data, for example, selecting a certain number of pixels from the entire image. Training on subsets eliminates dependence on data resolution, reduces memory consumption, and improves efficiency.
In particular, the recent development of neural implicit representation technology has made it a trend to widely distribute visual data by sharing implicit representation model weights. An example of this is the NeRF[9] for 3D representation. We believe that in the future, people will share various multimedia data contents that they capture as function formats, just as they currently share 2D images and videos online. In this context, we are also interested in the following research questions: (1) It is common to inject information into 2D images or 3D content for steganography or ownership recognition. But when people share information through implicit representation functions, can we achieve information hiding? (2) How to define steganographic capacity when implementing embedding operations in network with implicit representation?
@@
2 Related Work
2.1 Deep Steganography
@@
Deep steganography refers to a technique that employs deep neural networks to implement covert communication. These methods are mainly divided into two categories. One category is based on encoder-decoder networks. They utilize the powerful encoding ability of deep neural networks to encode the cover data and message into the stego data. Baluja[10] first proposed the use of deep neural networks as encoders to hide secret images into other images, and they also trained a decoder to recover the message. Zhu et al.[11] introduced a noise layer into this framework, allowing the decoder to achieve robust message recovery. Jing et al.[12] unified the encoder and decoder into a reversible network to achieve deep steganography. They also implemented multi-image embedding to achieve high-capacity steganography[13]. Xu et al. proposed a flow-based robust reversible image steganography framework[14]. Recently, a special method for steganography through fixed neural networks[15] has been proposed, which adds controllable noise to the carrier to avoid training the message extraction network. These methods require an existing natural and realistic cover to construct stego data.
@@
Another category of methods uses deep generative models, such as GANs, diffusion models, flow models, and autoregressive encoders to construct cover data, stego data or modification policies. Volkhonskiy et al.[16] first used GANs to generate an original cover image to counter steganalysis. Liu et al.[17] use a category-controlled GAN to establish a mapping relationship between the message and generated images to hide information. Similar to [16], Hu et al. establishe a mapping relationship between messages and noise images and then uses the GAN to generate a stego image directly[18]. Zhang et al. regarded the generator as an image sampler and extracted the message using a fixed message decoder called the Carden Grille[19]. Tang et al. used GANs to generate distortion cost functions to implement cover modification steganography[20, 21]. Chen et al. directly generates stego using variational autoencoders[22, 23] (VAE) and flow model-based generators[24][25]. Yang et al. proposed a new framework for generative steganography using autoregressive models, which uses the pixel CNN+ to generate hidden steganography[26]. Zhang et al.[27] also used autoregressive models to design steganography schemes and implemented pixel-level information hiding using Pixel CNN+ and arithmetic coding. Liu et al. decoupled images into structural and textural features using autoencoders, enabling stable feature extraction from structural features to improve the accuracy of secret message extraction[28]. Recently, there have also been schemes using diffusion models[29, 30] to construct steganography. Wei et al. embed the message using modified DC coefficients of the diffusion noise, and then a reverse diffusion model restores the noise to a natural image[31]. Finally, the message is extracted using a diffusion process and IDCT transform. Similar to Hu et al.[18], Kim et al.[32] establishes a mapping relationship between messages and noise images and then used a diffusion model to generate stego data. Yu et al. uses the image conversion ability of the diffusion model and the robustness of noise to adopt a stable diffusion conditional model to implement an image steganography[33].
@@
In order to achieve high-quality coding of cover data and decoding of messages, these deep learning-based steganography schemes usually require a lot of computing resources to obtain powerful encoders (generator) and decoders. These networks are usually trained on grid-like data such as images, making the size of these encoders and decoders directly related to the resolution of the underlying grid data. Except for [19] and [15], all the above methods need to transfer the message extractor with a large amount of data to the message receiver, which not only increases the communication burden but also increases the risk of exposing steganographic behavior. To solve this problem, Li et al.[7, 8] disguise a steganography network (called a secret DNN model) by ordinary deep neural networks (called stego DNN) with image classification or segmentation tasks. In [7], they disguised the entire secret message extractor by selecting and adjusting some of the filters to retain their functionality in the secret task. In [8], they proposed a gradient-based filter integration scheme to insert interference filters into important positions in the secret DNN model to form a stego DNN model. Although these two methods ensure the security of the message decoder, they still do not avoid the problem of high communication burden. These two methods use image processing behavior (classification or segmentation) to conceal steganography behavior. They only provide a concealment strategy for message extractors in steganographic schemes. In this paper, we present a complete steganographic framework for multimedia data and other data, where we represent the message as a continuous function and embed this function into another continuous function. This not only provides security for message extractors but also eliminates the need for additional transmission of stego data, as the stego function itself acts as both a stego and extractor. Additionally, the resolution of the neural implicit representation is decoupled from the training data, and implicit model representations typically have smaller data sizes, thus avoiding high-load communication.
2.2 Model Hiding
@@
With the widespread use of model data in the cloud, researchers have also paid extensive attention to information hiding in model data[34]. Model hiding can be mainly divided into two aspects: Neural Network Steganography and Neural Network Watermarking. Neural network watermarking can be classified into three types: white-box watermarking, black-box watermarking, and no-box watermarking. In white-box watermarking, The verifier can access the internal structure of the network, including the weights, to verify the copyright. Uchida et al. proposed a method to embed a watermark consisting of 0s and 1s into the weights of the network through regularization, and the watermark can be extracted using a specific algorithm to verify the copyright[23]. Black-box watermarking is suitable when the verifier cannot access the internal structure of the network and can only use remote API calls. A common approach is to select specific samples as triggers and fine-tune the network to fit these triggers for copyright verification[35]. No-box watermarking is mainly used for copyright verification of generative networks[36]. The strategy is to train the network to generate images that contain the watermark, and the verifier can directly verify the copyright from the generated images.
@@
Some scholars have proposed the model steganography where the model data is used as stego data for hidden information. Yang et al.[37] embedded secret data into the convolutional layers of a given neural network (the cover network) during the network training process. Matrix multiplication is used to encode the parameters of the convolutional layers for data extraction. Chen [38] used a deep neural network (DNN) to establish a probability density model of cover images and hide a secret image at a specific location of the learned distribution. They used SingGAN [39] to learn the distribution of patches in a cover image. During the patch distribution learning process, a deterministic mapping is fitted from a fixed set of noise maps (generated by embedding keys) to hide the secret image. Yang et al. proposed a multi-source data hiding scheme for neural networks, where multiple senders can simultaneously send different secret data to a receiver using the same neural network[40]. They achieved data embedding during the training process of the neural network, replacing post-training modifications to the network.
@@
2.3 Information Hiding For NeRF
Recently, a special type of neural implicit representation for 3D data called Neural Radiance Fields (NeRF)[9] has received extensive research attention. Li et al.[41] proposed StegaNeRF, a method to embed hidden information in NeRF rendering. They design an optimization framework that first trains a regular NeRF model and then performs a second training to achieve the rendering of the secret image while ensuring the original visual quality. Similarly, Luo et al. proposed an anti-distortion rendering scheme by replacing the original color representation in NeRF with a watermarked color representation, ensuring stable extraction of the watermark information from the 2D images rendered by NeRF.[42]. Chen et al.[43] utilized the synthesized novel views by NeRF as an important basis for copyright verification and trained a watermark verification model using a parameterized approach. Similar to [43], Dong et al. [44] applied this technique to steganography and developed a NeRf-based steganography scheme. These approaches treat the neural radiance field as a new-view image generator, embedding secret information in the images from the new views. Therefore, an additional decoder is still required for message extraction. In this paper, we embed the implicit representation function of the secret message into the implicit representation function of the cover data. We can directly extract the secret message from the stego function using the shared key between the sender and receiver, avoiding the need for the message extractor.
3 Representing Data As Functions
In this section, we will provide a comprehensive overview of INR, with a focus on using images as an illustrative example.
3.1 Representing an Image with Function
Suppose we have an image where $I[x;y]$ corresponds to the RGB value at pixel position $(x;y)$. We want to represent this image as a function $f:{{R}^{2}}\rightarrow{{R}^{3}}$, where $f(x;y)=(r;g;b)$ returns the RGB value at pixel position $(x;y)$. To achieve this, we parameterize the function $f_{\theta}$ with a multi-layer perceptron (MLP) with weighted parameters $\theta$, which is often referred to as an $INR$. We can then learn this representation by minimizing the following loss function[45] as described in Eq. (1).